# Conceptualization, Methodology, Writing, Visualization

### v1

@@ 3.3 Energy coordination and battery management
s.t. $\eta_{x}\sum_{t=t_{\mathrm{arr}}}^{t_{\mathrm{dep}}}p_{t}\cdot\Delta t=\mathrm{SoC}_{\mathrm{tar}}-\mathrm{SoC}_{\mathrm{arr}},$ (4)
@@ 3.4 Ancillary services
Due to the characteristics of electric vehicles, the EV battery can be considered both a power load and a distributed energy storage unit. Leveraging the energy storage function of the EV battery, EVs can provide power storage, power supply, and ancillary services to the distribution system when parked. With the modernization of power systems, integrating power and communication infrastructure based on smart grids enables energy exchange and information flow between EVs and the grid, which provides support for various services to enhance the reliability and sustainability of the power system. As a result, the V2G concept has emerged.
The ancillary services contain frequency regulation, spinning reserve, active power support, and reactive power compensation tan2016integration . Refs. rahman2022comprehensive ; tan2016integration ; guille2009conceptual ; kisacikoglu2010effects ; rotering2010optimal ; sortomme2011optimal ; ehsani2012vehicle pecifically discuss the implementation methods of V2G ancillary services. Ancillary services have the positive impact of maintaining the balance between the generation and load demand to ensure a stable and reliable power system. In addition, V2G services also help reduce emissions, increase profits, and provide additional income for electric vehicle owners. Meanwhile, using EVs to provide ancillary services presents several challenges. Beyond the issue of battery degradation from frequent bidirectional V2G operations, there is also the need for additional investment in bidirectional chargers.
@@ 3.4.1 Frequency regulation and Spinning reserve
Frequency regulation involves two modes: regulation up and regulation down pandit2024frequency ; almeida2015electric . These modes adjust the power levels of specific resources to fine-tune frequency by balancing supply and demand within a minute or less kempton2005vehicle . Specifically, when loads increase, EV charging rates can be reduced to achieve the regulation down mode in response to the increased load.
Spinning reserve provides synchronized generation capacity with fast response in a typical duration of minutes, generally within 10 minutes to compensate for the generation outage han2014practical ; bandyopadhyay2011aggregator . More specifically, EVs’ charging rates can be decreased by unidirectional V2G technologies to attain the additional spinning reserve.
@@ 3.4.2 Active power support and reactive power compensation
Reactive power compensation is a technique to provide voltage regulation and power factor correction tan2016integration , further increasing power system operating efficiency and reducing power loss, such as guille2009conceptual ; kisacikoglu2010effects . In conventional reactive power supply methods, a capacitive reactive power is needed for this function. Distribution generators and static Volt–Ampere Reactive (VAR) compensators are most commonly used. Based on the capacitive reactive power reserved in the DC-link capacitor of the EV bidirectional battery charger, the EV can supply reactive power compensation by controlling the AC/DC converter without any battery degradation. Methods of managing EV charging for mitigating negative impacts on the distribution system are presented in Figure 4.
Figure 4: Methods of managing EV charging for mitigating negative impacts on the distribution system.
@@ 4.1 State estimation
where $Z$ is the measurement vector of measurement, $X$ and $Y$ are vectors of state variables, $N$ is measurement noise, and $h$ is a function that relates state variables to measurements. Knowing the network topology is a prerequisite for evaluating the impact of EV charging on distribution networks taheri2019milp . However, due to the limited presence of meters and infrequently-calibrated line parameters, distribution grid operators have problems with accurate and updated real-time information on the grid topology, which causes the need for dynamic identification of network topology park2018exact . Studies alturki2018marginal ; taheri2019milp ; park2018exact have contributed to estimating grid topology according to data measured at smart meters and grid sensors sayed2022electric .
@@ 4.2 Real-time management system
The charging and discharging patterns of EVs are largely influenced by the unpredictable travel habits of their users. As EVs are integrated into the distribution system in an unpredictable way, significant uncertainties arise in the timing and location of grid loads. This unpredictability complicates load forecasting and increases the challenges of real-time grid monitoring and management. To address these issues, optimal charging strategies are often designed to maximize the use of renewable energy sources to meet the fluctuating demands of EVs. However, renewable energy generation is intermittent and variable, further aggravating uncertainties in real-time management. To address these issues, many studies have focused on handling the uncertainties of EV charging, renewable energy generation, real-time electricity price, and user behavior, where methods of stochastic optimization kaewdornhan2023real , scenario-based stochastic programming yan2023real ; eddahech2012real ; yao2016real , robust optimization zhao2015risk ; vaya2015self , and dynamic programming mohamed2013real ; you2016real ; binetti2015scalable ; liao2015dispatch have been widely applied. Though these approaches can effectively mitigate the impact of uncertainties, many face the challenge of computational dimensionality. For instance, capturing the characteristics of uncertainty requires numerous scenarios, which leads to a heavy computational burden. Even at a moderate scale, the stochastic programming for EV charging dispatching needs several hours to solve yan2023real . To address the problem of high computational intensity, decentralized and distributed methods have received considerable attention chokkalingam2017real ; csengor2020real ; li2023constrained . Compared to centralized schemes, decentralized approaches provide better scalability and real-time performance li2023constrained . li2023constrained proposes a decentralized framework and a dynamic energy boundary model of individual EVs to achieve discreteness of charging demands, which needs to train only a single agent model to enable local decision-making for multiple charging piles within a charging station. binetti2015scalable utilizes the scalable real-time greedy algorithm, zhou2020admm adopts alternating direction method of multipliers to schedule EVs in a decentralized method, and latifi2018agent designs a game theoretic model to keep both the local optimality and the global optimality for the aggregator. However, the aforementioned distributed methods have limited capability in addressing system-level objectives and overall cooperation requirements. Many studies focus on addressing the balance between mitigating the impact of uncertainties and high computational intensity. wu2017two proposes two control algorithms, SPLET and SAA_SPLET, to reduce the number of scenarios and alleviate the computational burden. li2023constrained introduces a dynamic energy boundary model for EVs to meet charging demand without penalty terms. wang2019reinforcement designes four feature functions to approximate the state-action function, reducing the dimension of the state space. The issue of EV charging and discharging dispatching is essentially a time arrangement for EV charging and discharging control while considering various uncertainties li2023constrained . The current control strategy will influence the system’s next state, which increases computational complexity and affects the convergence speed shen2021real . Therefore, formulating EV charging control issues as a Markov Decision Process (MDP) is a practical approach, where Model-free Deep Reinforcement Learning (DRL) techniques can be used as solutions for sequential decision-making problems zhang2018review to find optimal strategies in complex and uncertain environments.
@@
Due to the randomness of traffic conditions, user behavior, real-time electricity price, the arrival and departure time of EVs, as well as battery states and charging demands, are dynamic and challenging to predict accurately, which makes the management of EV charging challenging wan2018model . In recent years, many day-ahead scheduling methods zhao2015risk ; vaya2015self ; erdinc2014smart have been used to address this issue. However, they have limited capability in handling more complex real-time scenarios involving time-varying EV charging demand and electricity price. Real-time management systems with high efficiency and accuracy, along with appropriately designed DR mechanisms, are two key factors for DSOs (Distribution System Operators) in managing EV charging and discharging. The DR mechanism has already received extensive attention in akhavan2017managing ; ghavami2014nonlinear ; chekired2017decentralized , while there are still mismatches between day-ahead scheduling and real-time demand. Apart from the conventional solution of quickly starting up backup generators to provide a power supply, the EV battery system can also serve as a solution by treating EVs as flexible energy storage units capable of discharging electricity back into the grid. Besides, the EV V2B (Vehicle-to-Building) and SAB (Smart Aggregator-Based) systems can also be used as strategies to reduce operational costs wu2017two . Therefore, many studies have focused on enhancing the accuracy and efficiency of real-time management. For instance, in zhao2015risk , the state-action function is represented by a linear combination of feature functions to transform the decision in the time-varying action space into four time-invariant constants, solving the difficulty in representing the time-varying state and action space in EV charging. jiang2019real formulates the charging optimization problem as a cost minimization problem and proposes an improved binary grey wolf optimizer to improve the convergence speed and optimization accuracy. long2021efficient develops an ordinal optimization-based method to search for optimal charging strategies within seconds while providing a performance guarantee. yang2021dynamic introduces a stochastic dynamic simulation modeling framework for EV fast-charging real-time management, incorporating a multi-server queueing model and a multinomial logit station choice model based on charging prices, expected wait times, and detour distances. csengor2020real uses a linear programming method to propose an energy management model for electric vehicle parking lots based on real-time optimization to maximize load factor. yan2023real develops a real-time feedback integrated online algorithm based on Lyapunov optimization, which provides theoretical bounds for maximum charging delays.
Figure 5: Framework of real-time management system of EV charging.
@@
As large-scale EVs integrate into the power grid and the continuous development of real-time energy management systems, the scope of EV charging has expanded significantly. The coordination of bi-directional energy flows between EV charging systems and other systems, such as photovoltaic, thermal energy, and energy storage systems, has gained widespread attention. Many studies focus on real-time energy management, integrating EV charging with multiple interconnected systems. zhang2018dynamic proposes a dynamic cost optimization scheduling method based on real-time information of EV charging demand and PV generation to control the charging process of each EV in a parking lot without considering ESS. shen2021real introduced a Combined Heat and Power (CHP) system and proposed a real-time energy management algorithm for a microgrid based on the Lyapunov optimization technique to minimize the average cost of the microgrid. jiang2019real investigates the charging optimization for EVs in parking lots integrated with ESS and PV systems and proposes a real-time EV charging dispatching strategy based on an improved binary grey wolf optimizer. Figure 5 presents the framework of a real-time management system of EV charging. The Summary of the studies on the Real-time management system of EV charging is presented in Table 2.
@@
5 Conclusion and future outlook
This paper first discusses the impact of EV charging on the operation of power distribution systems from the perspective of power quality and grid stability. For the issue of real-time stable operation, the coping capacity of EV charging areas to grid faults and solutions to peak loads are reviewed. Further in-depth analysis is carried out to explore how to reduce the negative impact of EV charging on the distribution network and guide positive impacts, followed by a comprehensive summary of specific treatment methods. Considering that the state of charge of EV batteries and charging demand are difficult to predict, widespread random integration of EVs will introduce significant uncertainties into the real-time operation of the grid, increasing the complexity of real-time management. Therefore, this paper conducts an in-depth analysis of the real-time management of EV charging from the perspectives of monitoring and dispatching. First, real-time state estimation based on the topology of the distribution network is performed to assess the sensitivity of different locations in the grid to accommodate EV loads. Further strengthening of monitoring and management of weaknesses in the distribution network during operation is also necessary. Given the increasingly complex real-time scenarios of EV charging demand and price fluctuations, day-ahead scheduling methods have limited capabilities. While real-time scheduling methods usually face significant computational burdens when addressing uncertainties, balancing the impacts of uncertainties and computational intensity is crucial. Decentralized methods are effective for solving computationally intensive problems. However, they have limited capacity to address system-level goals and cooperative demands, leading to widespread attention to model-free deep reinforcement learning technologies. Moreover, it’s important to study highly efficient and accurate real-time management systems through improved algorithms, mechanisms, and connections with other systems. Regarding future research directions, enhancing the coping capacity of EV charging areas to grid faults is worthy of more attention. Additionally, the balance between the impacts of uncertainties and computational intensity needs to be further explored to improve the accuracy and efficiency of EV real-time management.

### v2

@@ 3.3 Energy coordination and battery management
s.t. $\eta_{p}\sum_{t=t_{\mathrm{arr}}}^{t_{\mathrm{dep}}}p_{t}\cdot\Delta t=\mathrm{SoC}_{\mathrm{tar}}-\mathrm{SoC}_{\mathrm{arr}},$ (4)
@@ 3.4 Ancillary services
The ancillary services contain frequency regulation, spinning reserve, active power support, and reactive power compensation tan2016integration . Refs. rahman2022comprehensive ; tan2016integration ; guille2009conceptual ; kisacikoglu2010effects ; rotering2010optimal ; sortomme2011optimal ; ehsani2012vehicle pecifically discuss the implementation methods of V2G ancillary services. Ancillary services can help maintain the balance between generations and loads to ensure stable and reliable power systems. In addition, V2G services also help reduce emissions, increase profits, and provide additional income for EV owners. Meanwhile, using EVs to provide ancillary services presents several challenges. Beyond the issue of battery degradation from frequent bidirectional V2G operations, there is also the need for additional investment in bidirectional chargers. Focusing on the issue of V2G, the automotive energy management system (EMS) deserves more attention due to the ability to coordinate the energy status of EVs under various conditions. For instance, zhang2025scenario proposes an integrated energy management strategy based on a multi-task deep reinforcement learning (DRL) algorithm, which dynamically adjusts the reserved SOC to optimize V2G participation while accounting for battery aging costs. Recognizing that reinforcement learning (RL) strategies often succeed in simulation but face challenges in real-world deployment, lei2025sim establishes an RL-based EMS development toolchain that leverages high-fidelity vehicle models for agent training. zhang2023integrated introduces a model-free, multi-state DRL algorithm for integrated thermal and energy management under cold climate conditions, which closes to dynamic programming fuel economy performance, with a margin of 93.7%.
@@ 3.4.1 Frequency regulation and Spinning reserve
The stability criterion of grid frequency is the power generation must match the load consumption, otherwise the frequency deviation from the criterion operating point will be caused peng2017dispatching . bu2019generic proposes a generic framework to tackle various frequency-related uncertainties and accommodate different system frequency response models. Given fast response time and low utilization time, EVs with V2G technology are suitable for providing frequency regulation services to the power system alfaverh2023optimal . Considering the three layers of frequency control, EVs can perform primary, secondary, and tertiary frequency regulation based on the generator droop characteristic simulation, area control error and economic dispatch, respectively xiao2013review . Frequency regulation involves two modes: regulation up and regulation down pandit2024frequency ; almeida2015electric . These modes adjust the power levels of specific resources to fine-tune frequency by balancing supply and demand within a minute or less kempton2005vehicle . Specifically, when loads increase, EV charging rates can be reduced to achieve the regulation down mode in response to the increased load. Spinning reserve provides synchronized generation capacity with fast response in a typical duration of minutes, generally within 10 minutes to compensate for the generation outage han2014practical ; bandyopadhyay2011aggregator . More specifically, EVs’ charging rates can be decreased by unidirectional V2G technologies to attain the additional spinning reserve.
The power level of EV discharging is small, thus EVs need to be aggregated and controlled by EV aggregator to perform frequency regulation. Specifically, when the EV aggregator receives the load frequency control (LFC) signal, it dispatches the signal into every single EV so that every regulation resource receive a control signal from the system operator. Dispatching control methods generally contains dispatching based on specifically designed rules, or a pro-rata basis by participation factors cai2022optimal . While rule-based method provides control flexibility, it is limited by complex control structures and sampling rate of the frequency regulation signal. Regulation signal intervals of 5–15 minutes karfopoulos2015distributed ; le2016pde cannot fully exploit the advantage of fast response. To overcome these limitations, cai2022optimal designs a dispatching control for LFC participation with a faster dispatching time-step to maximize the EV aggregator’s revenue, which is formulated as a nonlinear and nonconvex optimization problem and solved by a genetic algorithm. Simulation results demonstrate that the proposed approach enhances dispatching efficiency, increases regulation capacity, and yields a 6% revenue improvement. alfaverh2023optimal proposes an a DRL-based optimal V2G control strategy to simultaneously maximize the benefits of EV owners and aggregators while performing frequency regulation tasks, where a deep deterministic policy gradient agent is used to dynamically adjust the V2G power scheduling. jie2023contribution investigates the contribution of aggregated EV groups to system frequency regulation in a power grid with high PV penetration. Time-series simulations reveal that frequency deviation is reduced to 0% compared with the initial frequency of 3.5%.
@@ 3.4.2 Active power support and reactive power compensation
Reactive power compensation is a technique to provide voltage regulation and power factor correction tan2016integration , further increasing power system operating efficiency and reducing power loss, such as guille2009conceptual ; kisacikoglu2010effects . In conventional reactive power supply methods, a capacitive reactive power is needed for this function. Distribution generators and static volt–ampere reactive compensators are most commonly used. Based on the capacitive reactive power reserved in the DC-link capacitor of the EV bidirectional battery charger, the EV can supply reactive power compensation by controlling the AC/DC converter without any battery degradation. Methods of managing EV charging and discharging to mitigate negative impacts on distribution systems are presented in Figure 3.
Figure 3: Methods of EV management to mitigate negative impacts on distribution systems.
@@ 4.1 State estimation
where $Z$ is the measurement vector of measurement, $X$ and $Y$ are vectors of state variables, $N$ is measurement noise, and $h$ is a function that relates state variables to measurements. Knowing the network topology is a prerequisite for evaluating the impact of EV charging on distribution networks taheri2019milp . However, due to the limited deployment of meters and infrequently calibrated line parameters, distribution grid operators have problems with accurate and up-to-date real-time information park2018exact . Distribution-level phasor measurement unit (PMU) technology bu2022stability has a superior capability of monitoring, analyzing and controlling the real-time distribution systems, providing valid technical support for state estimation. Studies alturki2018marginal ; taheri2019milp ; park2018exact have contributed to estimating grid topology using data measured from smart meters and grid sensors sayed2022electric . Topology estimation in majority of the power distribution system is hindered by infrequently calibrated line parameters park2018exact and sparsely positioned monitoring devices deka2019topology , thus limiting the estimation of the rest of the system. To address this issue, taheri2019milp identifies the underlying grid topology by perturbing power injections and analyzing the corresponding voltage responses. This method follows a linear regression setup with unknown vector of line resistances. moffat2019unsupervised proposes a noise-robust technique to estimate effective impedances based on the reduced Laplacian form of the Kron reduced admittance matrix. Nevertheless, methods for dynamic topology updates remain limited, especially when considering EV integration. nie2016system proposes an effective system state estimation algorithm based on quasi-Newton method that integrates forecast charging load and predictable base power load. yao2025state proposes an EV aggregators rescheduling scheme for real-time congestion relief, based on a robust state estimation incorporating measured data from PMU and smart meters.
@@ 4.2 Real-time management system
The charging and discharging patterns of EVs are largely influenced by the unpredictable travel habits of their users. Real-time EV management system aims to dynamically generate charging schedules in response to time-varying charging demands and electricity prices wan2018model . As EVs are integrated into the distribution system in an unpredictable way, significant uncertainties arise in the timing and location of grid loads. This unpredictability complicates load forecasting and increases the challenges of real-time grid monitoring and management. To address these issues, optimal charging strategies are often designed to maximize the use of renewable energy sources to meet the fluctuating demands of EVs. However, renewable energy generation is intermittent and variable, further aggravating uncertainties in real-time management. To address these issues, many studies have focused on handling the uncertainties of EV charging, renewable energy generation, real-time electricity price, and user behavior, where methods of stochastic optimization kaewdornhan2023real , scenario-based stochastic programming yan2023real ; eddahech2012real ; yao2016real , robust optimization zhao2015risk ; vaya2015self , and dynamic programming mohamed2013real ; you2016real ; binetti2015scalable ; liao2015dispatch have been widely applied. Though these approaches can effectively mitigate the impact of uncertainties, many face the challenge of computational dimensionality. For instance, capturing the characteristics of uncertainty requires numerous scenarios, which leads to a heavy computational burden. Even at a moderate scale, the stochastic programming for EV charging dispatching needs several hours to solve yan2023real . Many studies focus on addressing the balance between mitigating the impact of uncertainties and high computational intensity. wu2017two proposes two control algorithms, SPLET and SAA_SPLET, to reduce the number of scenarios and alleviate the computational burden. li2023constrained introduces a dynamic energy boundary model for EVs to meet charging demand without penalty terms. wang2019reinforcement designes four feature functions to approximate the state-action function, reducing the dimension of the state space.
To address the problem of high computational intensity, decentralized and distributed methods have received considerable attention chokkalingam2017real ; csengor2020real ; li2023constrained . Compared to centralized schemes, decentralized approaches provide better scalability and real-time performance li2023constrained . rivera2014distributed proposes a distributed anytime algorithm for network utility maximization in real-time EV charging control. This algorithm can operate asynchronously without performance loss and achieve millisecond-level implementation, thereby improving robustness under fast dynamic conditions. The feasibility of results produced at each iteration ensures the response speed requirement. latifi2018agent develops a fully decentralized game-theoretic model to ensure both local optimality for individual EV and global optimality for the system aggregator. An arbitrary-private topology offers scalability while preserving customer privacy and resilience to node failures. Based on the simulations in a smart microgrid including one aggregator with ten customers, the proposed approach converges fast with one iteration per customer, and reduces grid payment compared with the unscheduled program. zhou2020admm designs a hierarchical algorithm based on the alternating direction method of multipliers to allow each individual EV to update its own charging strategy simultaneously, where a decentralized algorithm based on the gradient projection method is further employed to handle the non-separable load regulation term. Compared with the centralized method tested in a 5-feeder and a 12-feeder system with 350 EVs, the average time per outer iteration of the decentralized algorithm is 1.1s and the total computation time is 78s, which is much less than the 2057s required by the centralized method. binetti2015scalable proposes a scalable real-time greedy algorithm in a decentralized fashion, which departs from heavy computations and extensive bi-directional communications. Simulation results confirm its effectiveness under high EV penetration levels. li2023constrained proposes a decentralized framework based on recurrent deep deterministic policy gradient (RDDPG) and a dynamic energy boundary model of individual EV to achieve discreteness of charging demands, which needs to train only a single agent model to enable local decision-making for multiple charging piles within a charging station. A RDDPG-based 10-pile-model is tested with a cost reduction of 39.05% compared with the uncontrolled strategy of 50.42%. The proposed method demonstrates strong scalability and applicability to large-scale EVs without retraining. To validate this property, the 10-pile model is further extended to scenarios with 20, 30, 40, 50, and 100 piles, confirming the effectiveness under large-scale conditions.
However, the aforementioned distributed methods have limited capability in addressing system-level objectives and overall cooperation requirements. In this regard, the two-level hierarchical control framework has proven effective in improving scalability and separating concerns. The upper-level controller manages an aggregated representation of EVs and remains agnostic to individual charging behaviors, while the lower-level controller considers detailed models and the specific requirements of each EV engel2022hierarchical . However, the main drawback of this approach lies in its limited real-time responsiveness. To address this issue, hybrid strategies that combine long-term scheduling with short-term operational control have been applied latifi2018agent ; wu2017two , striking a balance between the computational burden of large-scale scenarios and the need for real-time decision-making. In essence, EV charging is a stochastic process influenced by multiple uncertainties that are difficult to quantify. Consequently, real-time management models that decouple from prediction models have attracted significant attention. Representative approaches include model-free reinforcement learning, Lyapunov optimization, and greedy algorithms. The greedy algorithm decomposes the offline problem into subproblems for each time slot, but it lacks guarantees of global optimality. Lyapunov optimization offers theoretical performance guarantees but is limited in handling highly complex state spaces. Model-free reinforcement learning requires substantial training but demonstrates strong adaptability in dynamic environments. The issue of EV charging dispatching is essentially a time arrangement for EV charging control while considering various uncertainties li2023constrained . The current control strategy will influence the system’s next state, which increases computational complexity and affects the convergence speed shen2021real . Therefore, formulating EV charging control issues as a Markov decision process is a practical approach, where model-free deep reinforcement learning (DRL) techniques can be used as solutions for sequential decision-making problems zhang2018review to find optimal strategies in complex and uncertain environments. wang2024two proposes a two-timescale Markov game to enable a model-free and decentralized EV charging control. An action-persistence multi-agent soft actor-critic algorithm is designed to ensure stationary by addressing the hybrid action space of remotely controlled switches and soft open points. bi2023real develops a DRL framework for real-time charging decisions, where an off-policy algorithm is designed for offline agent training. Since pre-trained models often fail when confronted with new EVs, changing grids, or evolving user behavior, zhang2025bi builds adaptive systems through lifelong or transfer learning, which enables rapid adaptation without retraining from scratch.
@@
Due to the randomness of traffic conditions, user behavior, real-time electricity price, the arrival and departure time of EVs, as well as battery states and charging demands, are dynamic and challenging to predict accurately, which makes the management of EV charging challenging wan2018model . In recent years, many day-ahead scheduling methods zhao2015risk ; vaya2015self ; erdinc2014smart have been used to address this issue. However, they have limited capability in handling more complex real-time scenarios involving time-varying EV charging demand and electricity price. Real-time management systems with high efficiency and accuracy, along with appropriately designed DR mechanisms, are two key factors for distribution system operators in managing EV charging and discharging. The DR mechanism has already received extensive attention in akhavan2017managing ; ghavami2014nonlinear ; chekired2017decentralized , while there are still mismatches between day-ahead scheduling and real-time demand. Apart from the conventional solution of quickly starting up backup generators to provide a power supply, the EV battery system can also serve as a solution by treating EVs as flexible energy storage units capable of discharging electricity back into the grid. Besides, the vehicle-to-building and smart aggregator-based systems can also be used as strategies to reduce operational costs wu2017two . Therefore, many studies have focused on enhancing the accuracy and efficiency of real-time management. For instance, in zhao2015risk , the state-action function is represented by a linear combination of feature functions to transform the decision in the time-varying action space into four time-invariant constants, solving the difficulty in representing the time-varying state and action space in EV charging. jiang2019real formulates the charging optimization problem as a cost minimization problem and proposes an improved binary grey wolf optimizer to improve the convergence speed and optimization accuracy. long2021efficient develops an ordinal optimization-based method to search for optimal charging strategies within seconds while providing a performance guarantee. yang2021dynamic introduces a stochastic dynamic simulation modeling framework for EV fast-charging real-time management, incorporating a multi-server queueing model and a multinomial logit station choice model based on charging prices, expected wait times, and detour distances. csengor2020real uses a linear programming method to propose an energy management model for electric vehicle parking lots based on real-time optimization to maximize load factor. yan2023real develops a real-time feedback integrated online algorithm based on Lyapunov optimization, which provides theoretical bounds for maximum charging delays.
Figure 4: Framework of real-time EV management system.
@@
As large-scale EVs integrate into the power grid and the continuous development of real-time energy management systems, the scope of EV charging has expanded significantly. The coordination of bi-directional energy flows between EV charging systems and other systems, such as photovoltaic, thermal energy, and energy storage systems, has gained widespread attention. Many studies focus on real-time energy management, integrating EV charging with multiple interconnected systems. zhang2018dynamic proposes a dynamic cost optimization scheduling method based on real-time information of EV charging demand and PV generation to control the charging process of each EV in a parking lot without considering ESS. shen2021real introduced a combined heat and power system and proposed a real-time energy management algorithm for a microgrid based on the Lyapunov optimization technique to minimize the average cost of the microgrid. jiang2019real investigates the charging optimization for EVs in parking lots integrated with ESS and PV systems and proposes a real-time EV charging dispatching strategy based on an improved binary grey wolf optimizer. Figure 4 presents the framework of a real-time management system of EV charging. The summary of the studies on the EV real-time management system is presented in Table 2.
@@
5 Conclusion and future outlooks
5.1 Summary of conclusion
This paper first discusses the impact of EV charging on the operation of power distribution systems from the perspective of power quality and grid stability. For the issue of real-time stable operation, the coping capacity of EV charging areas to grid faults and solutions to peak loads are reviewed. Further in-depth analysis is carried out to explore how to reduce the negative impact of EV charging on the distribution network and guide positive impacts, followed by a comprehensive summary of specific treatment methods. Considering that the state of charge of EV batteries and charging demand are difficult to predict, widespread random integration of EVs will introduce significant uncertainties into the real-time operation of the grid, increasing the complexity of real-time management. Therefore, this paper conducts an in-depth analysis of the real-time management of EV charging from the perspectives of monitoring and dispatching. First, real-time state estimation based on the topology of the distribution network is performed to assess the sensitivity of different locations in the grid to accommodate EV loads. Further strengthening of monitoring and management of weaknesses in the distribution network during operation is also necessary. Given the increasingly complex real-time scenarios of EV charging demand and price fluctuations, day-ahead scheduling methods have limited capabilities. While real-time scheduling methods usually face significant computational burdens when addressing uncertainties, balancing the impacts of uncertainties and computational intensity is crucial. Decentralized methods are effective for solving computationally intensive problems. However, they have limited capacity to address system-level goals and cooperative demands, leading to widespread attention to model-free deep reinforcement learning technologies. Moreover, it’s important to study highly efficient and accurate real-time management systems through improved algorithms, mechanisms, and connections with other systems.
5.2 Future research opportunities
The future research directions on V2G technologies are as follows:
  1. [1)]
  2. 1.
    EV-grid connections introduce new vulnerabilities, creating an urgent need for lightweight intrusion detection and privacy-preserving strategies.
  3. 2.
    Coordinating large and diverse EV fleets remains a major technical challenge, causing the need for effective real-time control approaches.
  4. 3.
    Current market mechanisms fail to adequately capture the value of V2G flexibility, emphasizing the necessity of developing new pricing and incentive models.
  5. 4.
    Future EV-grid systems will require the joint optimization of energy, traffic, and communication within a unified framework to ensure overall system efficiency and stability.
Additionally, enhancing the resilience of EV charging areas against grid faults deserves more attention. Dynamic state estimation considering EV integration requires further investigation to provide valid technical support for real-time operation systems. Furthermore, the trade-off between EV uncertainties and computational intensity should be further explored to improve the accuracy and efficiency of real-time EV management.